# A Note on the Phragmén-Lindelöf Theorem

### v2

@@ 3. Examples
which come from [patel_explicit_2022] (first three), [hiary2023explicit] (forth), and [BELLOTTI2024128249] or [leong2024explicitestimateslogarithmicderivative] (for the final). Additionally it is well known that for any $\eta>1$ we have
@@
One may verify numerically, the maximum modulus principal simplifying the check, that for $\sigma\in[1,2]$ and $t\in[0,3]$ that $\left|\frac{s-1}{s}\zeta(s)\right|\leq 1$ with the maximum being taken on at $s=1$. Also, for $\sigma\in[1,2]$ and $t\in[0,30]$ we have $\left|\frac{s-1}{s}\zeta(s)\right|\leq 2.1$ the maximum being near $1+27.7i$.
@@
Notice that for the cases of $G_{1}$, $G_{2}$, $G_{3}$ and $G_{5}$ the conditions for $|t|\leq 3$ is trivially verified as $\zeta(\eta)\geq\zeta(2)>1$. Similarly with $G_{4}$ with $|t|\leq 30$ and $\eta<1.5$ as $\zeta(1.5)>2.1$. To extend to $\eta\in[1.5,2]$ one can use that for $\sigma\in[1.5,2]$ and $t\in[0,30]$ we have $\left|\frac{\sigma+it-1}{\sigma+it}\zeta(s)\right|<\zeta(2)$. We thus conclude that for any $\eta\in[1,2]$ that for $\sigma={\rm Re}(s)\in[0,\eta]$ we have
@@
for all $t>t_{0}$. One verifies that in all cases we have
@@
[…]

### v3

@@ 4. Examples
We provide several examples illustrating how to apply our theorems to obtain bounds for the Riemann zeta function in contexts where the exponent on the log factor in the bound is decreasing in $\sigma$, constant in $\sigma$, or increasing in $\sigma$.
4.1. First Example
@@
which come from [undefe] (first three), [undefc] (fourth), and [undef] or [undefd] (for the final). Additionally it is well known that for any $\eta>1$ we have
@@
One may verify numerically, with the maximum modulus principle simplifying the check, that for $\sigma\in[1,2]$ and $t\in[0,3]$ that $\left|\frac{s-1}{s}\zeta(s)\right|\leq 1$ with the maximum being taken on at $s=1$. Also, for $\sigma\in[1,2]$ and $t\in[0,30]$ we have $\left|\frac{s-1}{s}\zeta(s)\right|\leq 2.1$ the maximum being near $1+27.7i$.
@@
Notice that for the cases of $G_{1}$, $G_{2}$, $G_{3}$ and $G_{5}$ the conditions for $|t|\leq 3$ are trivially verified as $\zeta(\eta)\geq\zeta(2)>1$. Similarly with $G_{4}$ with $|t|\leq 30$ and $\eta<1.5$ as $\zeta(1.5)>2.1$. To extend to $\eta\in[1.5,2]$ one can use that for $\sigma\in[1.5,2]$ and $t\in[0,30]$ we have $\left|\frac{\sigma+it-1}{\sigma+it}\zeta(s)\right|<\zeta(2)$. We thus conclude that for any $\eta\in[1,2]$ that for $\sigma={\rm Re}(s)\in[1,\eta]$ we have
@@
for all $t>t_{0}$ and $\sigma\in[1,\eta]$. One verifies that in all cases we have
@@
4.2. Second Example
To bound $|\zeta(s)|$ for $\Re(s)\in[5/7,1]$ we shall also need the bound
$|\zeta(5/7+it)|<1.546\log(t)t^{1/14}\qquad t\geq 3$
due to [undefj].
Now set $G(s)=\frac{1}{2}\left(\log(e+s)+\log(e+2-s)\right)$ and notice that for $t\geq 3$ and $0<\sigma<2$ we have
$\log(t)<\left|G(\sigma+it)\right|.$
Now set $H(s)=\zeta(s)\frac{s-1}{sG(s)}$. For $\sigma=5/7$ and $t\geq 3$ we have
$\left|H(\sigma+it)\right|\leq 1.546|\sigma+it|^{1/14}$
and for $\sigma=1$ and $t\geq 3$ we have
$\left|H(\sigma+it)\right|\leq 1.$
We apply our Theorem 1 to $H(s)$ to conclude that for $\sigma\in[5/7,1]$ we have
$|H(\sigma+it)|<1.546^{\frac{7}{2}(1-\sigma)}|\sigma+it|^{\frac{1}{4}(1-\sigma)}.$
It follows that for $\sigma\in[5/7,1]$ we have
$|\zeta(\sigma+it)|<1.546^{\frac{7}{2}(1-\sigma)}|\sigma+it|^{\frac{1}{4}(1-\sigma)}\left|\frac{\sigma+it}{\sigma+it-1}G(\sigma+it)\right|.$
Finally, for $t>10^{5}$ and $\sigma\in[5/7,1]$ we have
$\left|\frac{\sigma+it}{\sigma+it-1}G(\sigma+it)\right|<(1+10^{-10})\log|t|$
so that for $t>10^{5}$ and $\sigma\in[5/7,1]$ we have
$|\zeta(\sigma+it)|<(1+10^{-10})1.546^{\frac{7}{2}(1-\sigma)}|t|^{\frac{1}{4}(1-\sigma)}\log|t|.$ (6)
4.3. Third Example
To bound $|\zeta(s)|$ for $\Re(s)\in[1/2,5/7]$ we shall also need the bound
$|\zeta(\sigma+it)|\leq 66.7t^{\frac{27}{164}}$
due to [undeff].
We let $G(s)=\frac{1}{2}\left(\log(4e+s)+\log(4e+2-s)\right)$ so that $|G(\sigma+it)|$ is decreasing in $\sigma$ for $\sigma\in[1/2,5/7]$ and $t\geq 3$.
We thus have for $t\geq 3$ that
$|\zeta(1/2+it)|\leq 66.7t^{\frac{27}{164}}\leq 66.7(1/2+t)^{\frac{27}{164}}|G(1/2+it)|^{0}$
and
$|\zeta(5/7+it)|\leq 1.546|t|^{1/14}\log t\leq 1.546|5/7+it|^{1/14}|G(5/7+it)|.$
Now, since the minimum value of $1.546|\sigma+it|^{1/14}|G(\sigma+it)|$ and $66.7|\sigma+t|^{\frac{27}{164}}$ where $\sigma\in[1/2,5/7]$ and $t\in[0,3]$ are both larger than the maximum value of $|\zeta(\sigma+it)|$ in the same region we may apply Theorem 7 and conclude
$|\zeta(\sigma+it)|\leq\left(66.7|\sigma+it|^{\frac{27}{164}}\right)^{\frac{14}{3}(\frac{5}{7}-\sigma)}\left(1.546|\sigma+it|^{1/14}|G(\sigma+it)|\right)^{\frac{14}{3}(\sigma-\frac{1}{2})}.$
As before it is now straight forward to verify that for $t>10^{5}$ and $\sigma\in[1/2,5/7]$ we have
$|\zeta(\sigma+it)|\leq(1+10^{-9})66.7^{\frac{14}{3}(\frac{5}{7}-\sigma)}1.546^{\frac{14}{3}(\sigma-\frac{1}{2})}t^{\frac{47}{123}-\frac{107}{246}\sigma}(\log|t|)^{\frac{14}{3}(\sigma-\frac{1}{2})}.$ (7)